# The Generation Gap: Exploring Age Bias Underlying in the Value Systems of Large Language Models

### v2

@@ Abstract
In this paper, we explore the alignment of values in Large Language Models (LLMs) with specific age groups, leveraging data from the World Value Survey across thirteen categories. Through a diverse set of prompts tailored to ensure response robustness, we find a general inclination of LLM values towards younger demographics, especially in the US. Additionally, we explore the impact of incorporating age identity information in prompts and observe challenges in mitigating value discrepancies with different age cohorts. Our findings highlight the age bias in LLMs and provide insights for future work. Materials for our analysis will be available via anonymous.github.com
The Generation Gap: Exploring Age Bias Underlying in the Value Systems of Large Language Models
Siyang Liu1   Trish Maturi1   Bowen Yi1   Siqi Shen1   Rada Mihalcea1 1 The LIT Group, Department of Computer Science and Engineering, University of Michigan, Ann Arbor lsiyang@umich.edu, mihalcea@umich.edu
@@ 1 Introduction
Widely used Large Language Models (LLMs) should be reflective of all age groups Dwivedi et al. (2021); Wang et al. (2019); Hong et al. (2023). Age statistics estimate that by 2030, 44.8% of the US population will be over 45 years old Vespa et al. (2018), and one in six people worldwide will be aged 60 years or over World Health Organization (2022). Analyzing how the values (e.g, religious values) in LLMs align with different age groups can enhance our understanding of the experience that users of different ages have with an LLM. For instance, for an older group that may exhibit less inclination towards new technologies Czaja et al. (2006); Colley and Comber (2003), an LLM that embodies the values of a tech-savvy individual may lead to less empathetic interactions. Minimizing the value disparities between LLMs and the older population has the potential to lead to better communication between these demographics and the digital products they engage with.
Figure 1: Age-related bias in LLMs on thirteen human value categories. Human values in this figure refer in particular to the US groups. Trend coefficients (see calculation in Sec 3.3) were derived from the slope of the changing gap between LLM and human values as age increases. A positive trend coefficient signifies the widening gap observed from younger to older groups, thus indicating a model leaning towards younger age groups. Significant test is detailed in Appx G
In this paper, we investigate whether and which values in LLMs are more aligned with specific age groups. Specifically, by using the World Value Survey Haerpfer et al. (2020), we prompt various LLMs to elicit their values on thirteen categories, employing eight format variations in prompts for robust testing. We observe a general inclination of LLM values towards younger demographics, as shown in Fig 1. We also demonstrate the specific categories of value and example inquiries where LLMs exhibit such age preferences (See Sec 4).
Furthermore, we study the effect of adding age identity information when prompting LLMs. Specifically, we instruct LLMs to use an age and country identity before requesting their responses. Surprisingly, we find that adding age identity fails to eliminate the value discrepancies with targeted age groups on eight out of thirteen categories (see Fig 4), despite occasional success in specific instances (See Sec 5).
We advocate for increased awareness within the research community regarding the potential age bias inherent in LLMs, particularly concerning their predisposition towards certain values. We also emphasize the complexities involved in calibrating prompts to effectively address this bias.
@@ 2 Related Work
Due to the recent advancements in LLMs in manifesting human-level performance across various tasks Brown et al. (2020); Radford et al. (2019); Ouyang et al. (2022), there is a growing concern regarding the presence of social bias in these models Kasneci et al. (2023). Recent research has shown that LLMs exhibit “preferences” for certain demographic groups, such as White and female individuals Sun et al. (2023), and political inclination Santurkar et al. (2023); McGee (2023); Atari et al. (2023). Despite extensive scrutiny on LLM bias Santurkar et al. (2023); Sun et al. (2023), the age-related preferences of LLMs remain less explored. Previous work has mentioned age as one of multi-facets of bias in LLM performance Kamruzzaman et al. (2023); Haller et al. (2023); Draxler et al. (2023); Levy et al. (2024); Oketunji et al. (2023), while lacking a direct study on the age aspect. Recent research Duan et al. (2024) publishes an evaluation for well-known LLMs on age bias through 50 multi-choice questions; unlike it focuses on discriminatory narratives towards specific age groups, our investigation is running at an implicit level. We argue that understanding the underlying value systems is crucial, as the value discrepancies between users and LLMs can significantly impact their adoption of LLMs, even though the explicit discrimination is rectified, as exemplified in our discussion on technology attitudes in Sec 1.
Figure 2: Alignment rank of values of LLMs over different age groups in specific Countries. See results on more models and countries in Appx E and F . Rank 1 on an specific age group represents that this age group has the narrowest gap with LLM in values. A increasing monoticity indicates a closer alignment towards younger groups.
@@ 3.1 Human Data Acquisition
We derive human values utilizing the 7th wave of the World Values Survey (WVS) Haerpfer et al. (2020). The survey systematically probes 94k individuals globally on 13 categories, covering a range of social, political, economic, religious, and cultural values. See an introduction of WVS in Appx A. Each inquiry is a single-choice question. Responses are numeric, quantifying the inclination on the options, e.g., “1:Strongly agree, 2:Agree, 3:Disagree, 4:Strongly disagree”. Negative number is possible for coding exceptions such as “I don’t know”. To assess human values, we group the respondents by age group  [^1] Age groups are recorded as 18-24, 25-34, 35-44, 45-54, 55-64, and 65+ and country. Subsequently, we compute the average values for each age group and country to represent their respective cohorts, ignoring the invalid negative numbers.
@@ 3.2 Prompting
InstructGPT (GPT-3.5-turbo-instruct)
Mistral (mistral-7B-v0.1)
Vicuna (vicuna-7b-v1.5)
FLAN-T5 (flan-t5-xxl)
FLAN-UL2 (flan-ul2)
@@
We identify three key components for each inquiry in the survey: context, question ID&content, and options. To ensure robustness, we made several format variations for the prompt [^2] Despite adopting format variations, we were cautious to not include major changes as the content and structure of WVS were carefully designed by sociologists and professionals. (e.g., alter wordings and change order of components), as previous research Shu et al. (2023); Röttger et al. (2024); Beck et al. (2023) uncovered inconsistent performance in LLMs after receiving a minor prompt variation. Eventually, we build a set of eight distinct prompts per inquiry. Please see prompt design details in Tab 3 in Appendix. Through a careful analysis on the prompt responses (Appx B), We observe unstableness of LLM’s responses to prompt variations. However, multiple prompt trials assists with achieving a convergence point. On 95.5% of questions, more than half of the eight prompts led to responses centered on the same choice or adjacent options, and thus we believe it is acceptable to consider the average of the outcomes across the eight prompt variations as the LLM’s final responses to WVS. In addition, due to the instability of LLMs in following instructions, we encountered seven types of unexpected reply and present our coping methods for each, as summarized in Tab 4. In the process of averaging responses, we ignore the invalid negative numbers, as we did in calculating human values. For reproducing our work, parameter setting and prompting details are reported in Appendix D.
@@ 3.3 Measures
By collecting 372 value vectors that represent people across 62 countries and 6 age groups, along with a value vector for the LLM to compare, we utilize principle component analysis (PCA) Tipping and Bishop (1999) on totally 373 value vectors for representation learning. We acquire value representations for all groups with the dimensionality of three. Our consideration of using PCA is added in Appx C.
[…]
Let $i$ be the index of age group in [18-24, 25-34, 35-44, 45-54, 55-64, 65+] and the value representation for the $i$th age group be $x_{c,i},y_{c,i},z_{c,i}]$. We derive three metrics below for our further analyses:
@@
Trend Coefficient, the slope of the value gap between LLM and humans across six age groups. $\alpha$ is the slope we would like to fit by linear regression.
@@ 4 Aligning with Which Age on Which Values?
As shown in Fig 1, we observe a general inclination of popular LLMs favoring the values of younger demographics in the US on different value categories, indicated by the trend coefficient. Fig 2 exemplifies the bias for LMMs across six age groups in several countries. Due to the limited paper pages, results on other LLMs and countries can be found in Appx E and F. Significant testing procedure is available in Appx G. We observe that in the US and China, as countries of large population, the models tend to have a higher alignment rank on younger groups on the most categories, despite few exceptions (e.g., happiness and well-being). However, in Ethiopia and Nigeria (Tab 8), the inclination is less evident. We leave this phenomenon for future study.
@@
Case Study.
In Fig 3, we show two representative prompts and their responses from ChatGPT and human groups, to illustrate sample values where ChatGPT exhibits a clear bias toward a specific age group.
@@ 5 The Effect of Adding Identity in Prompts
To analyze if adding age identity in the prompt helps to align values of LLM with the targeted age groups, we adjust our prompts by adding a sentence like “Suppose you are from [country] and your age is between [lowerbound] and [upperbound].” at the beginning of the required component of the original prompt and get responses that corresponds with six age groups.
@@
We illustrate the change of Euclidean distance between values of LLM and different age groups after adding identity information. As is presented in Fig 4, in eight out of thirteen categories (No.1,2,4,5,7,8,9,12) no improvement is observed.
@@
We also showcase a successful calibration example for a question about the source of acquiring information in Fig 5. The value pyramid illustrates LLMs’ responses for different age ranges compares to the answers from the U.S. population. When age is factored into the LLM prompt, the LLM’s views are more aligned with the U.S. population of that respective age group, as it reports higher frequency using radio news for the older group.
@@
6 Recommendations for Future Work
We have observed that simply including an age in prompts fails to eliminate the value disparity for the targeted age groups. Out of the thirteen categories inquired upon, eight have shown no improvement. To this end, we recommend a careful data curation during pretraining. Doing so involves a deliberate and thoughtful selection of data sources that are diverse and representative of various age groups. By doing so, we can ensure that the model’s training material reflects a wide range of perspectives and experiences, thereby reducing biases and disparities in the model’s responses. We also recommend a consideration of human feedback optimization (e.g., RLHF). Through this iterative process, LLMs can learn to generate responses that fit better with the needs of different age groups. These strategies help mitigate the value disparities associated with targeted age groups, enhancing the LLM’s abilities to be more equitable and inclusive.
@@ 7 Conclusion
In this paper, we investigated the alignment of values in LLMs with specific age groups using data from the World Value Survey. Our findings suggest a general inclination of LLM values towards younger demographics. Our study contributes to raising attention to the potential age bias in LLMs and advocate continued efforts from the community to address this issue. Moving forward, efforts to calibrate value inclinations in LLMs should consider the complexities involved in prompting engineering and strive for equitable representation across diverse age cohorts.
@@
There are several limitations in our paper. Firstly, due to the time and cost, we were not able to try more sophisticated prompts for the age alignment, which may effectively eliminate the value disparity with targeted age groups. Secondly, our analysis relies on the questionnaire of WVS. However, their question design is not perfectly tailored for characterizing age discrepancies, which limits the depth of sights we could get from analysis. Finally, the range of LLMs in our analysis could be expanded.
@@
Several ethical considerations have been included thorough our projects. Firstly, the acquisition of WVS data is under the permission of data publisher. Secondly, we carefully present our data analysis results with an academic honesty. This project is under a collaboration, we well-acknowledge the work of each contributor and ensure a transparent and ethical process thorough the whole collaboration. Finally, we leverage the ability of AI-assistants to help with improving paper writing while we guarantee the originality of paper content and have reviewed the paper by every word.

### v3

# The Generation Gap: Exploring Age Bias in the Value Systems of Large Language Models

@@ Abstract
We explore the alignment of values in Large Language Models (LLMs) with specific age groups, leveraging data from the World Value Survey across thirteen categories. Through a diverse set of prompts tailored to ensure response robustness, we find a general inclination of LLM values towards younger demographics, especially when compared to the US population. Although a general inclination can be observed, we also found that this inclination toward younger groups can be different across different value categories. Additionally, we explore the impact of incorporating age identity information in prompts and observe challenges in mitigating value discrepancies with different age cohorts. Our findings highlight the age bias in LLMs and provide insights for future work. Materials for our analysis are available at https://github.com/MichiganNLP/Age-Bias-In-LLMs
The Generation Gap: Exploring Age Bias in the Value Systems
of Large Language Models
Siyang Liu  Trisha Maturi  Bowen Yi  Siqi Shen  Rada Mihalcea The LIT Group, Department of Computer Science and Engineering, University of Michigan, Ann Arbor lsiyang@umich.edu, mihalcea@umich.edu
Figure 1: Age-related bias in LLMs on thirteen human value categories. Human values in this figure refer in particular to the US groups. Trend coefficients (see calculation in Sec 3.3) were derived from the slope of the changing gap between LLM and human values as age increases. A positive trend coefficient signifies the widening gap observed from younger to older groups, thus indicating a model leaning towards younger age groups. The significance test is detailed in Appx F.
@@ 1 Introduction
Widely used Large Language Models (LLMs) should be reflective of all age groups Dwivedi et al. (2021); Wang et al. (2019); Hong et al. (2023). Age statistics estimate that by 2030, 44.8% of the US population will be over 45 years old Vespa et al. (2018), and one in six people worldwide will be aged 60 years or over World Health Organization (2022). Analyzing how the values (e.g., religious values) in LLMs align with different age groups can enhance our understanding of the experience that users of different ages have with an LLM. For instance, for an older group that may exhibit less inclination towards new technologies Czaja et al. (2006); Colley and Comber (2003), an LLM that embodies the values of a tech-savvy individual may lead to less empathetic interactions. Minimizing the value disparities between LLMs and the older population has the potential to lead to better communication between these demographics and the digital products they engage with.
In this paper, we investigate whether and which values in LLMs are more aligned with specific age groups. Specifically, by using the World Value Survey Haerpfer et al. (2020), we prompt various LLMs to elicit their values on thirteen categories, employing eight format variations in prompts for robust testing. We observe a general inclination of LLM values towards younger demographics, as shown in Fig 1. We also demonstrate the specific categories of value and example inquiries where LLMs exhibit such age preferences (See Sec 4). Furthermore, we study the effect of adding age identity information when prompting LLMs. Specifically, we instruct LLMs to use an age and country identity before requesting their responses. Surprisingly, we find that adding age identity fails to eliminate the value discrepancies with targeted age groups on eight out of thirteen categories (see Fig 4), despite occasional success in specific instances (See Sec 5). We advocate for increased awareness within the research community regarding the potential age bias inherent in LLMs, particularly concerning their predisposition towards certain values. We also emphasize the complexities involved in calibrating prompts to effectively address this bias.
@@ 2 Related Work
Due to the rapid advancements in LLMs across various tasks Brown et al. (2020); Ouyang et al. (2022), there is a growing concern regarding the presence of social bias in these models Kasneci et al. (2023). Recent research has shown that LLMs exhibit “preferences” for certain demographic groups, such as White and female individuals Sun et al. (2023), and political inclination McGee (2023); Atari et al. (2023). However, the age-related preferences of LLMs remain less explored. Prior work has mentioned age as one of multi-facets of bias in LLM performance Kamruzzaman et al. (2023); Haller et al. (2023); Draxler et al. (2023); Levy et al. (2024); Oketunji et al. (2023) while lacking a direct study on the age aspect. Recent research Duan et al. (2024) publishes an evaluation for well-known LLMs on age bias through 50 multi-choice questions; unlike it focuses on discriminatory narratives towards specific age groups, our investigation is running at an implicit level. We argue that understanding the underlying value systems is crucial, as the value discrepancies between users and LLMs can significantly impact their adoption of LLMs, even when the explicit discrimination is rectified, as exemplified in technology attitudes discussed in Sec 1.
@@ 3.1 Human Data Acquisition
We derive human values utilizing a well-established survey dataset, the 7th wave of the World Values Survey (WVS) Haerpfer et al. (2020). The survey systematically probes 94k individuals globally on 13 categories, covering a range of social, political, economic, religious, and cultural values. See more about WVS in Appx A. Each inquiry is a single-choice question. Responses are numeric, quantifying the inclination on the options, e.g., “1:Strongly agree, 2:Agree, 3:Disagree, 4:Strongly disagree". Negative number is possible for coding exceptions such as “I don’t know". To assess human values, we group the respondents by age group 11118-24, 25-34, 35-44, 45-54, 55-64, and 65+ and country. Subsequently, we compute the average values for each age group and country to represent their respective cohorts, ignoring the invalid negative numbers.
@@ 3.2 Prompting
[…]
*[Uncaptioned image] 0
[…]
*[Uncaptioned image] 0
[…]
*[Uncaptioned image] 0
[…]
*[Uncaptioned image] 0
[…]
*[Uncaptioned image] 0
[…]
*[Uncaptioned image] 0
[…]
*[Uncaptioned image] 0
[…]
*[Uncaptioned image] 0
[…]
*[Uncaptioned image] 0
[…]
*[Uncaptioned image] 0
[…]
*[Uncaptioned image] 0
[…]
*[Uncaptioned image] 0
[…]
*[Uncaptioned image] 0
[…]
*[Uncaptioned image] 0
[…]
*[Uncaptioned image] 0
[…]
*[Uncaptioned image] 0
[…]
*[Uncaptioned image] 0
[…]
*[Uncaptioned image] 0
@@
Figure 2: Alignment rank of values of LLMs over different age groups in specific Countries. See results on more models and countries in Appendix D and E . Rank 1 on a specific age group means that this age group has the narrowest gap with LLM in values. An increasing monoticity indicates a closer alignment towards younger groups.
@@
We identify three key components for each inquiry in the survey: context, question ID&content, and options. To ensure robustness, we made several format variations for the prompt [^2] Despite adopting format variations, we were cautious to not include major changes as the content and structure of WVS were carefully designed by sociologists and professionals. (e.g., alter wordings and change order of components), as previous research Shu et al. (2023); Röttger et al. (2024); Beck et al. (2023) uncovered inconsistent performance in LLMs after receiving a minor prompt variation. Eventually, we build a set of eight distinct prompts per inquiry. Please see prompt design details in Tab 8. Through a careful analysis of the prompt responses (Appx B), we observe the unstableness of LLM’s responses to prompt variations. However, multiple prompt trials assist with achieving a convergence point. On 95.5% of questions, more than half of the eight prompts led to responses centered on the same choice or adjacent options, and thus we believe it is acceptable to consider the average of the outcomes across the eight prompt variations as the LLM’s final responses to WVS. In addition, due to the instability of LLMs in following instructions, we summarize seven types of unexpected replies and present our coping methods for each in Tab 3. In the process of averaging responses, we ignore the invalid negative numbers, as we did in calculating human values. For reproducing our work, prompting details are reported in Appx C.
@@ 3.3 Measures
By collecting 372 value vectors that represent people across 62 countries and 6 age groups, along with a value vector for the LLM to compare, we perform min-max normalization, normal standardization, and then conduct principle component analysis (PCA) Tipping and Bishop (1999) on a total of 373 value vectors for representation learning. We acquire value representations for all groups with the dimensionality of three. Our consideration of using PCA is in Appx G.1.
[$x_{c},y_{c},z_{c}]=PCA\_transform([r_{1},r_{2},...r_{n_{c}}])$
Let $i$ be the index of age group in [18-24, 25-34, 35-44, 45-54, 55-64, 65+] and the value representation for the $i$th age group be $[x_{c,i},y_{c,i},z_{c,i}]$. We derive three metrics below for our further analyses:
@@
Trend Coefficient, the slope of the value gap between LLM and humans across six age groups. Let $\alpha_{c}^{*}$ be the optimal coefficient to fit the linear relation:
@@
Our reasons for these measure designs are detailed in the Appx G.
@@ 4 Aligning with Which Age on Which Values?
Fig 2 exemplifies the bias for LLMs across six age groups in several countries. Due to the limited paper pages, results on other LLMs and countries can be found in Appx D and E. As it is not intuitive to see a bias towards younger people in these decoupled results, we summarize the performance of all LLMs in the US, as shown in Fig 1. Then we observe a general inclination of popular LLMs favoring the values of younger demographics in the US on different value categories, indicated by the trend coefficient. Significance testing procedure is available in Appx F. We observe that in the US and China, as countries with large populations, the models tend to have a higher alignment rank on younger groups on most categories, despite few exceptions (e.g., happiness and well-being). However, in Ethiopia and Nigeria (Tab 15), the inclination is less evident. We leave this phenomenon for future study.
Case Study.
In Fig 3, we show two representative prompts and their responses from ChatGPT and human groups, to exemplify values where ChatGPT displays a clear inclination toward a specific age group. Note LLM values can be far away from all human age groups, as depicted in the second sub-figure. We discuss this point in Appx G.2.
@@ 5 The Effect of Adding Identity in Prompts
To analyze if adding age identity in the prompt helps to align values of LLM with the targeted age groups, we adjust our prompts by adding a sentence like “Suppose you are from [country] and your age is between [lowerbound] and [upperbound].” at the beginning of the required component of the original prompt and get responses that correspond with six age groups.
@@
We illustrate the change of Euclidean distance between values of LLM and different age groups after adding identity information. As is presented in Fig 4, in eight out of thirteen categories (No.1,2,4,5,7,8,11,12) no improvement is observed.
@@
We also showcase a successful calibration example for a question about the source of acquiring information in Fig 5. The value pyramid illustrates LLMs’ responses for different age ranges compared to the answers from the U.S. population. When age is factored into the LLM prompt, the LLM’s views are more aligned with the U.S. population of that respective age group, as it reports higher frequency using radio news for the older group.
@@
6 Further Discussion on the Age Bias Observed in LLMs
In this study, we have shown how LLMs are not representative of the value systems of older adults. Although further validation is necessary for a solid conclusion, we believe there may be several potential harms arising from this bias:
  • Older adults tend to place greater trust in established organizations, particularly when it comes to security concerns (as illustrated in Fig 1). An LLM unaware of these differences may pose greater risks to older users, who may be less prepared to identify misinformation from what appears to be a credible source (e.g., LLM itself). This could amplify the harm caused by LLM-generated hallucinations when letting LLMs serve aged people.
  • LLMs may offer less empathetic interactions to older adults by failing to account for their traditional beliefs, leading to less respectful exchanges.
  • For older adults, who are often less inclined towards new technologies, interacting with LLMs embodying the values of tech-savvy users could further alienate them. As shown in Fig 3, many older adults still rely on the radio for news, while younger people predominantly use the internet.
7 Suggestions on Age-aware Alignment for Future Work
Although we have shown that LLMs are not representative of the value systems of older adults, our study is not intended to promote a naive copy of the values of different age groups to achieve alignment. Simplistically applying statistical knowledge of the values of a particular age group might reinforce stereotypes rather than promote genuine alignment. For example, consider whether LLMs should adopt the value that the older generation is less tech-savvy and thus develop the stereotype that an older user would primarily obtain news from the radio rather than social media. However, as illustrated in Fig 6, while fewer older adults rely on social media for information, a significant portion still does. Therefore, LLMs must be aware of statistical discrepancies but should avoid brute-force applying statistics to any individual, as a brute-force application often only considers the mean instead of other qualities, such as variance, outliers, and so on. Thus, to facilitate a true age-aware alignment, we recommend researchers to rely on the following rules of thumb:
  • Avoid naively applying statistical knowledge of the values of a particular age group, as this can reinforce stereotypes instead of promoting genuine alignment.
  • Develop strategies that promote true age-sensitive interactions, emphasizing age-aware helpfulness and harmlessness, grounded in an understanding of value discrepancies across generations.
Figure 6: People’s preference on obtaining information from social media across different age groups in the US population
Achieving age-aware alignment requires LLMs to be sensitive to value differences across age groups and to build on these insights to offer helpful and harmless responses. For example, when engaging with older users, instead of brute-force assuming they are lagging behind new technology, a well-aligned system should keep tracking their understanding of the ongoing topics, offering more detailed explanations and minimizing the use of neologisms only when confusion arises. To achieve such age-sensitive interactions, exploring an effective feedback-acquiring method during interactions that complies with the real age-tailored connotation of helpfulness and harmlessness is meaningful. Although challenging, we believe this is a vital direction for future research.
@@ 8 Conclusion
In this paper, we investigated the alignment of values in LLMs with specific age groups using data from the World Value Survey. Our findings suggest a general inclination of LLM values towards younger demographics. Our study contributes to raising attention to the potential age bias in LLMs and advocates continued efforts from the community to address this issue. Moving forward, efforts to calibrate value inclinations in LLMs should consider the complexities involved in prompt engineering and strive for equitable representation across diverse age groups.
@@
There are several limitations in our paper. Firstly, Fig 3 may raise questions concerning the importance of any trends in light of LLM values not resembling any age group of humans. We conjecture that due to the nature of Human Preference Optimization Rafailov et al. (2024); Ouyang et al. (2022), LLMs develop extreme preferences (e.g., manifest an extreme atheist). The resulting LLMs will thus be unlike the subtler preferences of humans. Our study does not focus on the absolute difference between LLMs and humans, but instead emphasizes the inclination, as we have explained in Appendix G.2. However, future work is needed to reflect on the current process of Human Preference Optimization, especially on whether it will be problematic or acceptable if we over-align LLMs with human preference. Secondly, due to time and cost considerations, we were not able to try more sophisticated prompts for age alignment, which may effectively eliminate the value disparity with targeted age groups. Finally, our analysis relies on the questionnaire of WVS. However, their question design is not perfectly tailored for characterizing age discrepancies, which limits the depth of sight we could get from analysis.
@@
Several ethical considerations have been included through our projects. Firstly, the acquisition of WVS data is under the permission of the data publisher. Secondly, we carefully present our data analysis results with academic honesty. This project is under a collaboration, we well-acknowledge the work of each contributor and ensure a transparent and ethical process throughout the whole collaboration. Finally, we leverage the ability of AI assistants to help with improving paper writing while we guarantee the originality of paper content and have reviewed the paper by every word.
Acknowledgements
We thank the anonymous reviewers for their constructive feedback, and the members of the Language and Information Technologies lab at the University of Michigan for the insightful discussions during the early stage of the project. We thank Trenton Chang for his insightful questions in AI snippet activity at the University of Michigan. This project was partially funded by a National Science Foundation award (#2306372) and a grant from OpenAI. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation or OpenAI.